LabelEnergy

A-3.000×10-19J

B-5.000×10-19J

C-8.000×10-19J

D-1.400×10-18J

Label each of the following photons as ones that can or cannot be absorbed by this atom without ionizing the atom. The value stated is the energy of the photon.
Can be absorbed Cannot be absorbed  1.20×10-18J
Can be absorbed Cannot be absorbed  1.00×10-19J
Can be absorbed Cannot be absorbed  6.00×10-19J
Can be absorbed Cannot be absorbed  7.00×10-19J
Can be absorbed Cannot be absorbed  3.00×10-19J
Answer #1

the photons that can be absorbed from given energy levels are

1) (5 - 3) x 10^-19 j = 2.0 x 10^-19 J

2) (8 - 3) x 10^-19 j = 5.0 x 10^-19 J

3)  (14 - 3) x 10^-19 j = 11.0 x 10^-19 J = 1.1 x 10^-18 J

4)  (8 - 5) x 10^-19 j = 3.0 x 10^-19 J

5)  (14 - 5) x 10^-19 j = 9.0 x 10^-19 J

6)  (14 - 8) x 10^-19 j = 6.0 x 10^-19 J

hence from given photons only  6.00 x 10^-19J and  3.00 x 10^-19J can be absorbed.